### v2

# Gibbs randomness-compression proposition: An efficient deep learning

Abstract
A proposition that connects randomness and compression is put forward via Gibbs entropy over set of measurement vectors associated with a compression process. The proposition states that a lossy compression process is equivalent to directed randomness that preserves information content. The proposition originated from the observed behaviour in newly proposed Dual Tomographic Compression (DTC) compress-train framework. This is akin to tomographic reconstruction of layer weight matrices via building compressed sensed projections, via so-called weight rays. This tomographic approach is applied to previous and next layers in a dual fashion, that triggers neuronal-level pruning. This novel model compress-train scheme appears in iterative fashion and acts as a smart neural architecture search, The experiments demonstrated the utility of this dual-tomography producing state-of-the-art performance with efficient compression during training, accelerating and supporting lottery ticket hypothesis. However, random compress-train iterations having similar performance demonstrated the connection between randomness and compression from statistical physics perspective, we formulated the so-called Gibbs randomness-compression proposition, signifying randomness-compression relationship via Gibbs entropy. Practically, the DTC framework provides a promising approach for massively energy- and resource-efficient deep learning training.
Index Terms:
Compression algorithms, Compressive sensing, Information entropy, Deep learning architecture
I Introduction
Connections between randomness and data compression are noted earlier within the algorithmic information framework of Solomonoff-Kolmogorov-Chaitin [1, 2, 3] (SKC). In summary these works aim at reaching a definition of randomness from model selection perspective [4], i.e., algorithmic entropy or a complexity measure. Striking connections to algorithmic entropy and Gibbs entropy are also explored from fundamental conceptual point of view [5, 6, 7]. Statistical mechanics connections and information are established between Gibbs and Boltzmann entropies [8, 9, 10, 11]. However, relationship among compression process, entropy, information and randomness is always made quite generic in SKC framework. This manifestation was by design quite generic without strict mathematical condition, i.e., aimed at finding the shortest algorithm or explanation that can generate the random sequence. This is similar in entropy production studies [12].
In this work, we propose a mathematical condition among these concepts in a more formal manner inspired by newly proposed deep learning compress-train framework. We describe a technique called inverse compressed-sensing in dual fashion, Dual Tomographic Compression (DTC) compress-train for deep learning in an iterative manner as an example of Gibbs randomness-compression proposition. Such interplay between physics and deep machine learning is now well-established [13, 14, 15, 16, 17, 18].
II Background
Our purpose is establishing a mathematical relationship between randomness and complexity, whereby complexity appears as a compression process. In this direction, we leverage a couple of distinct research areas, randomness and complexity, compressed sensing and neural network architecture compression.
The definition of randomness [19] and quality of generating random numbers [20] appears to be connected subjects within the complexity, compressing data [6], empirically [21] and SKC algorithmic randomness context. This framework commonly known as Kolmogorov Complexity (KC), algorithmic complexity or algorithmic randomness and its applications are vast [22, 23]. KC or SKC framework, makes a generic statement about information content of sequence $s$ of binary digits of length $N$, $K(s)$, defined to be the size of the smallest program in bits that generates $s$, the sequence. The connection to compression is implicit here and its mathematical definition to compression process appears as only a single compression step. It is hinted that there is a positive correlation between compression ratio and randomness [22]. Moreover, the interest of randomness in this manner also explored from biological evolution stand point [24], which is more relevant in our case due to proposed train-compress framework having mutation-like process.
Compressed sensing (CS) framework has shown huge leap in statistical signal processing and information theory as it shows one can reconstruct an unknown signal with much less measurements under sparsity constraints [25, 26, 27, 28, 29]. CS involves random sampling and it provides an excellent tool to investigate the relationship between compression and randomness.
Reducing size of neural networks has attracted research interest since their inception, such as earlier works on Hopfiled networks [30] in statistical physics community. This pioneering work identifies clipping as removing existing connections and pruning re-learning, i.e., fine-tuning after clipping. Similar clipping approach first applied in computer science by LeCun et al. [31], fun title of brain damage is used, reflecting the performance degradation due to clipping, hinting at magnitude pruning, so called small saliency that small weights don’t affect the performance that much, also recently for semi-supervised learning [32, 33]. With the advent of deep learning [34, 35] and generative models producing ever increasing size large deep learning architectures due to attention [36] and diffusion [37] mechanisms, necessitates producing much smaller deep learning models for both training and inference efficiency, as energy consumption and ethical use is raised as the most important aspect in industry and academic practice [38] and sparse model size reduction post-model training became important [39], for dynamic cases as well [40, 41]. Lottery ticket hypothesis [42], provides a deeper perspective in compressing or finding smaller deep learning architectures, via magnitude pruning, i.e., brain damage, stating that subnetworks are possible that give similar performance as the initial deep architecture due to initialization chance, i.e., a lottery ticket. Applying the lottery ticket approach, iteratively pruning during training, leads to train-compress procedure, a compression process, which is the main example tool for our quest for finding relationship between randomness and compression. Such process of train-compress as a meta algorithm has been already proposed [43, 44], before lottery tickets were proposed. Effectiveness of random only pruning also reported recently [45]. Using compressed sensing in pruning recurrent networks is also explored, however it was only on to reduce size of independently appearing network components not on connections [46]. Using deep learning within compressed sensing framework is also explored recently in a different angle [47, 48].
The contributions of the paper are presented in the following sections Inverse Compressed Sensing, Dual Tomographic Compression and the main proposition, Gibbs randomness-compression proposition. These are supported by in-depth state-of-the-art (SoTA) numerical experiments on canonical vision deep learning test system.
Figure 1: DTC accuracy over different sparsity levels in train-compress iterative pruning.
Figure 2: DTC accuracy over different sparsity levels in train-compress iterative pruning. Inset shows a magnified view of higher sparsity levels.
III Inverse Compressed Sensing
We apply compressed sensing on the learned weights. However, as weights are known, instead of measuring the signal (weights), we do have it. This implies so called inverse compressed sensing, not to be confused with inverse problem nature of the original compressed sensing framework. Inverse here implies, we construct an hypothetical measurement process from known signal, and try to reconstruct the signals sparse version via conventional compressed sensing, i.e.,so called weight rays. In case of deep learning layers, total weights associated to neuron’s input and output will be used instead of single CS, dual nature of tomographic process. Essentially, we found a sparse projection of the learned weights, which might be sparser due to compressed sensing framework. In the notation, we will use weights, instead of calling a signal vector [28]. Feedforward fully-connected weights are assumed, however generalizing to other layer types should be obvious.
Given, learned weights vector $w\in\mathbb{R}^{m}$, that is flatten from fully connected layer matrix, representing all neurons state at a given learning time. Our purpose is to find sparse projection of $w$, $w_{s}\in\mathbb{R}^{n}$ where by $n$ is defined via sparsity level $s$, $n=m\cdot s$. Within CS framework, we define two matrices, one measurement matrix $\Psi\in\mathbb{R}^{mxm}$, here discrete cosine transform matrix and randomization matrix, here normal Gaussian, $\Phi\in\mathbb{R}^{nxm}$. Then CS matrix reads $\Theta=\Phi\cdot\Psi\in\mathbb{R}^{nxm}$. The inverse bit as we discusses, comes to play when we generate a hypothetical measurement vector, $y=\Phi\cdot w$ as a measurement vector. Then we minimize the following unconstrained version of CS procedure, to reconstruct the learned weights, so called weight ray $w_{r}$,
$\min_{w_{r}}\left\lVert\Theta w_{r}-y\right\rVert^{2}+\lambda\left\lVert w_{r}\right\rVert_{1}$
Obtaining different $w_{r}$ given different sparsity levels $s$ corresponds to neuronal level pruning in our canonical example, if magnitudes ranked in some manner. However, such inverse compressed sensing procedure can be applied to any compression process whereby hypothetical measurement could be generated.
Figure 3: Weight-rays, total of reconstructed weight over different sparsity levels.
IV Dual Tomographic Compression: Neuronal-level pruning
Compressed sensing approach proposed in an inverse fashion can be utilized for neuronal-level pruning. This can be achieved by applying inverse compressed sensing in a dual fashion on a layer with $n$ neurons. Weight vector can be obtained as follows: a fully-connected previous layer with $p$ neurons, weight matrix $W_{p}\in\mathbb{R}^{pxn}$ and the next layer with $q$ neurons, weight matrix $W_{q}\in\mathbb{R}^{nxq}$, are used to sum over rows, resulting per neuron values in obtaining weight rays $w_{p}\in\mathbb{R}^{n}$ and $w_{q}\in\mathbb{R}^{n}$. These weight vectors are reconstructed via inverse compressed sensing, in dual fashion:
$\min_{w_{r}^{p}}\left\lVert\Theta w_{r}^{p}-y_{p}\right\rVert^{2}+\lambda\left\lVert w_{r}^{p}\right\rVert_{1}$
$\min_{w_{r}^{q}}\left\lVert\Theta w_{r}^{q}-y_{q}\right\rVert^{2}+\lambda\left\lVert w_{r}^{q}\right\rVert_{1}$
Then the reconstructed sparse projections summed, $w_{r}^{p}+w_{r}^{q}$ are used to clip neurons where their weight rays’ sum is lower than the target sparsity. This is achieved via clipping at the quantile value corresponding to the sparsity level, between $0.0$ and $1.0$, from no sparsity to higher sparsity values, respectively. This compression is applied at each learning cycle, $m$ batches, forming the so-called train-compress cycle. After clipping, we retain the weights from previous learning cycle and continue to train further at this new reduced number of neurons as a pruning step, i.e., fine-tuning. This entire iterative pruning procedure is called Dual Tomographic Compression (DTC).
V Gibbs randomness-compression proposition
The proposition is inspired by DTC’s formulation in the previous section. Given a compression process of $n$ steps on a vector $w$, it progresses with measurement vectors $y_{i}$, $i=1,...,n$ in an iterative fashion with a random mechanism ${\bf R}$, as each step corresponds to a compression ratio $s_{i}$ and a fitness value, i.e., performance metric $f_{i}$. The sizes of measurement vectors are monotonically decreasing as the compression ratio compares to the initial signal. A normalized distribution of measurements is formed and its Gibbs entropy is computed as $G_{i}$, i.e., from normalized histogram bins of $k$, where $G_{i}=-\Sigma_{k}p_{k}log_{2}p_{k}$.
The Gibbs randomness-compression proposition states that there is an almost perfect relationship between the two functions $(s_{i},f_{i})$ and $(s_{i},G_{i})$. This proposition establishes that directed randomness, because of its iterative nature, within the compression process is related to compression performance. As compression with preserved fitness is related to the level of complexity, the proposition establishes a relationship between randomness and complexity. In other words, a lossy compression process is equivalent to directed randomness that preserves information content, as evidenced by the retention of performance or fitness.
Figure 4: Previous layer Gibbs entropy of computed measurements within DTC over compression process.
VI Numerical results
We have embedded DTC into a canonical vision task, classifying MNIST dataset [49] via softmax output. Basic deep learning models are implemented with hidden layer size of 512 neurons as starting point for train-compress iterative pruning. A fixed seed Xavier initialization [50] is used for reproducibility. We apply dual compressed sensing at every 200 batches of size 512, triggering inverse compressed sensing procedure using SCS algorithm [51] for minimizing for weight rays, for simplicity we didn’t use bias terms. We utilize pytorch [52] for deep learning and cvxpy [53] for DTC steps.
TABLE I: Test accuracy for different pruning approaches over different sparsity levels in iterative pruning.
Sparsity Magnitude Prune Random Prune DTC Prune
0.1016 0.9638 0.9637 0.9663
0.1914 0.9675 0.9669 0.9687
0.3477 0.9735 0.9723 0.9728
0.4141 0.9745 0.9736 0.9735
0.5254 0.9740 0.9747 0.9732
0.6172 0.9730 0.9748 0.9732
0.7520 0.9718 0.9719 0.9705
0.8594 0.9560 0.9588 0.9609
0.9121 0.9324 0.9399 0.9463
0.9219 0.9234 0.9341 0.9423
0.9297 0.9144 0.9292 0.9371
0.9512 0.8703 0.9157 0.9153
0.9707 0.8293 0.8834 0.8717
0.9785 0.8138 0.8673 0.8500
0.9824 0.8060 0.8589 0.8391
We record the test accuracy of the reduced architecture after each train-compress cycle for different pruning approaches: magnitude only, random and DTC based magnitute pruning. As shown in Table 1, Figure 1 and Figure 2. DTC and random pruning perform superior. We observe that DTC shows superior performance against random pruning after 85 percent sparsity. For very high sparsity levels after 95 percent, randomness shows a bit better performance but this might be due to very small network that has quite fluctuating results. Depending on requirements very good performance is also retained even after 95 percent of the neurons are pruned.
Figure 5: Next layer Gibbs entropy of computed measurements within DTC over compression process.
Total weight rays in Figure 3 are shown to be more pronounced for very high sparsity levels. This signifies, attenuation of learned weight as model gets much smaller to compensate the performance loss. This is a striking result as DTC sparsify the weights via inverse compressed sensing.
We also reported Gibbs entropy over train-compress iterative pruning in Figures 4 and 5. As complexity reduces, Gibbs entropy over measurement vectors also decreases. Pearson correlations between these and performance reduction are computed for DTC and random pruning. They read 0.9174 and 0.9412 respectively, empirically supporting our proposition.
Further study on the effect of frequency of dual tomographic procedure during train-compress iterations, variation of the resulting performances as we fixed the initialization for exact reproducibility, as estimating uncertainty reliably requires very large repeats, more optimizations on the penalty term of inverse compressed sensing procedure and the effect of binning on the Gibbs entropy on the measurement vectors appears to be good avenues to explore.
VII Conclusion
Contributions of the work appear to be quite stimulating for both theory and practice in three folds 1) we have formed mathematically more explicit relationship between directed randomness and complexity reduction compared to algorithmic randomness framework of SKC and demonstrated this empirically on our train-compress framework on the canonical deep learning task of classifying handwritten images. 2) formulation of an inverse compressed sensing framework, whereby measurements are generated rather than being a primary process 3) a new train-compress iterative pruning approach at neuronal level for deep learning is shown as a promising new approach for immensely energy efficient deep learning, supporting the lottery tickets hypothesis as well.
Acknowledgment
We are grateful to Y.Süzen for her kind support of this work and her role in securing private funding. Especially her encouragement to make the prototype codes public domain in order to contribute to energy-efficient AI for a better world.
Supplement
Two Jupyter notebooks are provided for prototype implementations and reproducing the results presented in this work, named $dtcp\_analysis.ipynb$ and $dtcp\_generate\_data.ipynb$