### v3

# Gibbs randomness-compression proposition: An efficient deep learning

Abstract
A proposition that connects randomness and compression is put forward via Gibbs entropy over set of measurement vectors associated with a compression process. The proposition states that a lossy compression process is equivalent to directed randomness that preserves information content. The proposition originated from the observed behavior in newly proposed Dual Tomographic Compression (DTC) compress-train framework. This is akin to tomographic reconstruction of layer weight matrices via building compressed sensed projections, via so-called weight rays. This tomographic approach is applied to previous and next layers in a dual fashion, that triggers neuronal-level pruning. This novel model compress-train scheme appears in iterative fashion and acts as a smart neural architecture search: also called compression aware training. The experiments demonstrated the utility of this dual-tomography during training: method accelerates and supports lottery ticket hypothesis. However, random compress-train iterations having similar performance demonstrated the connection between randomness and compression from statistical physics perspective, we formulated the so-called Gibbs randomness-compression proposition, signifying randomness-compression relationship via Gibbs entropy. The proposition is supported with the experimental evidence, resulting in very high correlation between learning performance vs. Gibbs entropy over compression ratios. Practically, the DTC framework provides a promising approach for massively energy- and resource-efficient deep learning training.
keywords:
Compression algorithms, Compressive sensing, Information entropy, Deep learning architecture
journal:
1 Introduction
Connections between randomness and data compression are noted earlier within the algorithmic information framework of Solomonoff-Kolmogorov-Chaitin [solomonoff64, kolmogorov65, chaitin66] (SKC). In summary these works aim at reaching a definition of randomness from model selection perspective [chaitin75], i.e., algorithmic entropy or a complexity measure. Striking connections to algorithmic entropy and Gibbs entropy are also explored from fundamental conceptual point of view [bennett82, zurek89, machta99a]. Statistical mechanics connections and information are established between Gibbs and Boltzmann entropies [jaynes65a, jaynes57a, buonsante16, rondoni00]. However, relationship among compression process, entropy, information and randomness is always made quite generic in SKC framework. This manifestation was by design quite generic without strict mathematical condition, i.e., aimed at finding the shortest algorithm or explanation that can generate the random sequence. This is similar in entropy production studies [suzen25a]. In this work, we propose a mathematical condition among these concepts in a more formal manner inspired by newly proposed deep learning compress-train framework. We describe a technique called inverse compressed-sensing in dual fashion, Dual Tomographic Compression (DTC) compress-train for deep learning in an iterative manner as an example of Gibbs randomness-compression proposition. Such interplay between physics and deep machine learning is now well-established [sherrington, hopfield82, ackley85, suzen22c, suzen2023h, suzen24].
2 Background
In order to establish a mathematical relationship between randomness and complexity, where the concept of complexity appears as a compression process, a concrete cases are indispensable. In this direction, we leverage a couple of distinct research areas, randomness and complexity, compressed sensing and neural network architecture compression, as a foundational background of the work.
The definition of randomness [compagner91] and quality of generating random numbers [oneill14] appears to be connected subjects within the complexity, compressing data [zurek89], empirically [chang10] and SKC algorithmic randomness context. This framework commonly known as Kolmogorov Complexity (KC), algorithmic complexity or algorithmic randomness and its applications are vast [li08, zurek91]. KC or SKC framework, makes a generic statement about information content of sequence $s$ of binary digits of length $N$, $K(s)$, defined to be the size of the smallest program in bits that generates $s$, the sequence. The connection to compression is implicit here and its mathematical definition to compression process appears as only a single compression step. It is hinted that there is a positive correlation between compression ratio and randomness [li08]. Moreover, the interest of randomness in this manner also explored from biological evolution stand point [wagner12], which is more relevant in our case due to proposed train-compress framework having mutation-like process.
Compressed sensing (CS) framework has shown huge leap in statistical signal processing and information theory as it shows one can reconstruct an unknown signal with much less measurements under sparsity constraints [donoho06a, candes06, candes06tao, baraniuk07, eldar12a]. CS involves random sampling and it provides an excellent tool to investigate the relationship between compression and randomness.
Reducing size of neural networks has attracted research interest since their inception, such as earlier works on Hopfield networks [janowsky89] in statistical physics community. This pioneering work identifies clipping as removing existing connections and pruning re-learning, i.e., fine-tuning after clipping. Similar clipping approach first applied in computer science by LeCun et al. [lecun89], fun title of brain damage is used, reflecting the performance degradation due to clipping, hinting at magnitude pruning, so called small saliency that small weights don’t affect the performance that much, also recently for semi-supervised learning [shwartz2024compress, shani2025tokens]. With the advent of deep learning [lecun15a, schmidhuber15a] and generative models producing ever increasing size large deep learning architectures due to attention [vaswani17] and diffusion [sohl15] mechanisms, necessitates producing much smaller deep learning models for both training and inference efficiency, as energy consumption and ethical use is raised as the most important aspect in industry and academic practice [menghani23, zhu25] and sparse model size reduction post-model training became important [hoefler2021sparsity], for dynamic cases as well [han2021dynamic, tatarnikova2025optimization]. Lottery ticket hypothesis [frankle18], provides a deeper perspective in compressing or finding smaller deep learning architectures, via magnitude pruning, i.e., brain damage, stating that subnetworks are possible that give similar performance as the initial deep architecture due to initialization chance, i.e., a lottery ticket. Applying the lottery ticket approach, iteratively pruning during training, leads to train-compress procedure, a compression process, which is the main example tool for our quest for finding relationship between randomness and compression. Such process of train-compress (or compression aware training) as a meta algorithm has been already proposed [han15, carreira18, zimmer25], before lottery tickets were proposed. Effectiveness of random only pruning also reported recently [gadhikar23]. Using compressed sensing in pruning recurrent networks is also explored, however it was only on to reduce size of independently appearing network components not on connections [zhen21a]. Using deep learning within compressed sensing framework is also explored recently in a different angle [wu2019deep, machidon2023deep]. The contributions of the paper are presented in the following sections Inverse Compressed Sensing, Dual Tomographic Compression and the main proposition, Gibbs randomness-compression proposition. These are supported by in-depth learning task on canonical vision deep learning test system as a baseline experimental evidence.
Figure 1: DTC accuracy over different sparsity levels in train-compress iterative pruning.
Figure 2: DTC accuracy over different sparsity levels in train-compress iterative pruning. Inset shows a magnified view of higher sparsity levels.
3 Inverse Compressed Sensing
We apply compressed sensing on the learned weights. However, as weights are known, instead of measuring the signal (weights), we do have it. This implies so called inverse compressed sensing, not to be confused with inverse problem nature of the original compressed sensing framework. Inverse here implies, we construct an hypothetical measurement process from known signal, and try to reconstruct the signals sparse version via conventional compressed sensing, i.e.,so called weight rays. In case of deep learning layers, total weights associated to neuron’s input and output will be used instead of single CS, dual nature of tomographic process. Essentially, we found a sparse projection of the learned weights, which might be sparser due to compressed sensing framework. In the notation, we will use weights, instead of calling a signal vector [baraniuk07].
Given, learned weights vector $w\in\mathbb{R}^{m}$, that is flatten from any layer matrix without any dependency on the layer type, representing all neurons state at a given learning time. Our purpose is to find sparse projection of $w$, $w_{s}\in\mathbb{R}^{n}$ where by $n$ is defined via sparsity level $s$, $n=m\cdot s$. Within CS framework, we define two matrices, one measurement matrix $\Psi\in\mathbb{R}^{mxm}$, here discrete cosine transform matrix and randomization matrix, here normal Gaussian, $\Phi\in\mathbb{R}^{nxm}$. Then CS matrix reads $\Theta=\Phi\cdot\Psi\in\mathbb{R}^{nxm}$. The inverse bit as we discusses, comes to play when we generate a hypothetical measurement vector, $y=\Phi\cdot w$ as a measurement vector. Then we minimize the following unconstrained version of CS procedure, to reconstruct the learned weights, so called weight ray $w_{r}$,
$\min_{w_{r}}\left\lVert\Theta w_{r}-y\right\rVert^{2}+\lambda\left\lVert w_{r}\right\rVert_{1}$
Obtaining different $w_{r}$ given different sparsity levels $s$ corresponds to neuronal level pruning in our canonical example, if magnitudes ranked in some manner. However, such inverse compressed sensing procedure can be applied to any compression process whereby hypothetical measurement could be generated.
Figure 3: Weight-rays, total of reconstructed weight over different sparsity levels.
4 Dual Tomographic Compression: Neuronal-level pruning
Compressed sensing approach proposed in an inverse fashion can be utilized for neuronal-level pruning. This can be achieved by applying inverse compressed sensing in a dual fashion on a layer with $n$ neurons. Weight vector can be obtained as follows: a fully-connected previous layer with $p$ neurons, weight matrix $W_{p}\in\mathbb{R}^{pxn}$ and the next layer with $q$ neurons, weight matrix $W_{q}\in\mathbb{R}^{nxq}$, are used to sum over rows, resulting per neuron values in obtaining weight rays $w_{p}\in\mathbb{R}^{n}$ and $w_{q}\in\mathbb{R}^{n}$. These weight vectors are reconstructed via inverse compressed sensing, in dual fashion:
$\min_{w_{r}^{p}}\left\lVert\Theta w_{r}^{p}-y_{p}\right\rVert^{2}+\lambda\left\lVert w_{r}^{p}\right\rVert_{1}$
$\min_{w_{r}^{q}}\left\lVert\Theta w_{r}^{q}-y_{q}\right\rVert^{2}+\lambda\left\lVert w_{r}^{q}\right\rVert_{1}$
Then the reconstructed sparse projections summed, $w_{r}^{p}+w_{r}^{q}$ are used to clip neurons where their weight rays’ sum is lower than the target sparsity. This is achieved via clipping at the quantile value corresponding to the sparsity level, between $0.0$ and $1.0$, from no sparsity to higher sparsity values, respectively. This compression is applied at each learning cycle, $m$ batches, forming the so-called train-compress cycle. After clipping, we retain the weights from previous learning cycle and continue to train further at this new reduced number of neurons as a pruning step, i.e., fine-tuning. This entire iterative pruning procedure is called Dual Tomographic Compression (DTC).
4.1 Interactions between consecutive reconstructions
Here, we take one of the simplest interactions as $w_{r}^{p}+w_{r}^{q}$ for clipping neurons in a ranked manner. However, as this is a framework more complex interactions could be developed further but our purpose here is to establish the DTC’s utility as a framework.
4.2 Layer types and propagation over layers
We express the framework for fully-connected layers but DTC doesn’t depends on the layer types within the neural network. Only requirement is mapping layer into weight rays, i.e., mapping incoming and outgoing connections into vectors. This can be achieved trivially by flattening any tensor representations of the layers to a one dimensional vector for DTC procedure, such as convolution or transformer architectures.
DTC should be applied to each hidden-layer back-to back, so called in forward propagating the DTC. Not to be confused with backpropagation. This is a simple propagation of the DTC procedure. In the case of multiple hidden-layers, DTC procedure becomes easier as incoming connections would already pruned due to clipped neurons from the previous layer.
We present this train-compress cycles in Algorithm 1 as an overall summary. Algorithm description aimed at presenting a representation of the framework with its logical components, rather than implementation details, in order to convey the core idea.
Input: NN number of learn-compress cycles, $L$ hidden-layers with $j$ as index, $n_{j}$ number of neurons at layer $j$, weight rays $w_{j}^{p}$ and $w_{j}^{q}$, $s_{i}$ target rate of prunning (decreasing monotonically)
Output: $L$ hidden layers with reduced size
Data: Training set $X_{tr}$, testing set $X_{te}$
/* Learn-compress cycles: Within batch learning, smaller network at each batches */
1 for i=1,,Ni=1,...,N do
2   Train the network on the given batch(s) taken from $X_{tr}$
 /* Layer-wise compression: DTC propagating over layers. */
3 for j=1,,Lj=1,...,L do
4      Inverse Compressed sensing:
5      Compute measurement matrices $\Theta$ and measurements vector $y$.
6      Dual $\ell_{1}$ minimization, find weights rays $w_{j}^{p}$ and $w_{j}^{q}$.
7      Clipping the layer (neurons) based on quantile at $s_{i}$ percent.
8      Ranking is achieved with using $w_{j}^{p}+w_{j}^{q}$ in order.
9      Compress the layer: Update layers with kept neurons, updating $n_{j}$
10
 /* For L >1 take mean of all Gibbs entropies */
11   Compute Gibbs entropy $G_{i}(s_{i})$ over measurement vectors $y_{j}$.
12   Compute the performance $f_{i}(s_{i})$ of the compressed network on the test set $X_{te}$
Algorithm 1 An algorithmic summary of Dual Tomographic Compression (DTC) based train-compress cycles. This is a representation of the general approach in the framework.
4.3 Computational Complexity
An assessment of the computational complexity of applying DTC follows. The primary compute overhead due to DTC are optimization cycles. This burden is compensated by the train-compress approach. Because in this framework we don’t need to complete entire training before applying the compression, recall compression aware training. We quantify this by the number of optimization calls over cycles against if we were to run full training and compression. $M$ cycles of compression and $N$ batches for learning and compressions, number of Layers $L$.
  • 1.
    DTC with train-compress
    Order of $N\cdot L$ calls to optimizers both learning and compression in batch learning. Note that in this case $N=M$ as we embed the compression within the training loop.
  • 2.
    Full train compress without DTC
    Order of $N\cdot M$ calls to optimizers for learning and assuming compression wouldn’t need an optimizer call.
We have deduced that, even though DTC brings some optimization load, computational complexity is compensated in a manner that, DTC framework is less compute heavy compare to full training and compress frameworks comparatively as $M$ compression cycles are usually much larger than the number of layers $L$ due to gradual compression, i.e., because compressing in large reductions in small steps are quite unlikely.
5 Gibbs randomness-compression proposition
The proposition is inspired by DTC’s formulation in the previous section. Given a compression process of $n$ steps on a vector $w$, it progresses with measurement vectors $y_{i}$, i=1,,ni=1,...,n in an iterative fashion with a random mechanism $\bf R$}, as each step corresponds to a compression ratio $s_{i}$ and a fitness value, i.e., performance metric $f_{i}$. The sizes of measurement vectors are monotonically decreasing as the compression ratio compares to the initial signal. A normalized distribution of measurements is formed and its Gibbs entropy is computed as $G_{i}$, i.e., from normalized histogram bins of $k$, where $G_{i}=-\Sigma_{k}p_{k}log_{2}p_{k}$.
The Gibbs randomness-compression proposition states that there is an almost perfect relationship between the two functions $(s_{i},f_{i})$ and $(s_{i},G_{i})$, denoted by $f_{i}(s_{i})$ and $G_{i}(s_{i})$. This proposition establishes that directed randomness, because of its iterative nature, within the compression process is related to compression performance. As compression with preserved fitness is related to the level of complexity, the proposition establishes a relationship between randomness and complexity. In other words, a lossy compression process is equivalent to directed randomness that preserves information content, as evidenced by the retention of performance or fitness.
5.1 Analysis of the proposition
More formally, we express the idea in terms of relationship between two comonotonic functions, being highly correlated.
Theorem 5.1 (Gibbs-randomness proposition).
Given sequential lossy compression process consists of $M$ consecutive compression cycle. At each compression cycle reduces the data size $s_{i}$ percent, representing model size, with relative to uncompressed size, $i=1,..,M$, where $s_{0}=1.0$. At each compression cycle, a measure vector ${\bf y}_{i}$ due to a randomized algorithm with associated Gibbs entropy $G_{i}(s_{i})$ and a corresponding learning performance $f_{i}(s_{i})$ after the compression cycle i.e., performance on the reduced model, are generated. If learning performance degradation and entropy reduction due to compression kept small in consecutive cycles $|f_{i}-f_{i-1}|<\epsilon_{i}$ and $|G_{i}-G_{i-1}|<\delta_{i}$ during entire learn-compress cycles, then functions $f_{i}$ and $G_{i}$ are highly correlated.
Proof.
A logical proof follows, functions $f_{i}(s_{i})$ and $G_{i}(s_{i})$ are non-increasing monotonically (comonotonic) due to quasi-static decreases $\epsilon_{i}$ and $\delta_{i}$ locally, while mean differences $\Delta$ are close to zero, $\Delta(\epsilon_{i})=\Delta(\delta_{i})\rightarrow 0$ as $|s_{i}-s_{i-1}|\rightarrow 0$. These functions are comoving over compress-learn cycle (functions evolve synchronously), hence are highly correlated. ∎
We known that functions $f_{i}(s_{i})$ and $G_{i}(s_{i})$ are mechanistically connected due to fact that measurement vector ${\bf y}_{i}$ is defined on the network weights, and learned weights directly effect the network learning performance. Because of random process, here Dual Tomographic Compression (DTC) generates ${\bf y}_{i}$, hence the theorem establishes the connection between randomness and compression.
The theorem is expressed for a single hidden-layer but it is readily applicable to entire neural network by simplify repeating it to all hidden-layers back to back at a given compress-train cycle. It can be applied to any layer type by mapping the weight structure to measurement vector ${\bf y}_{i}$. This is demonstrated in the Algorithm 1 in detail.
Figure 4: Previous layer Gibbs entropy of computed measurements within DTC over compression process.
6 Experiments and discussions
We have embedded DTC into a canonical vision task, classifying MNIST dataset [lecun02] via softmax output. A deep learning model is implemented with hidden layer size of 512 neurons as starting point for train-compress iterative pruning.
We apply dual compressed sensing at every 200 batches of size 512, triggering inverse compressed sensing procedure using SCS algorithm [odono16] for minimizing for weight rays, for simplicity we didn’t use bias terms. The selection of 200 batches were made such that performance is increasing sufficiently, even it is gradual, and limited by the trade-off between compute budget and compressing in a meaningful rate.
Hyperparameter and related details include Adam optimizer’s learning rate of $0.001$, ReLUReLU activations, fixed Xavier initialization [kumar17a] , batch normalization for the baseline network architecture with fully connected layers. We utilize PyTorch [paszke17] for deep learning and cvxpy [diamond16] for DTC steps.
In the conducted experiments, two more different clipping approach is implement for fair comparison. Namely random and magnitude pruning. Random clipping is just randomly drops neurons at the specified sparsity levels. In the case of magnitude pruning we simply rank the highest weight inputs to given neurons. In total we run three compress-train cycles to compare the test performances, as summarized in Table 1.
Table 1: Test accuracy for different pruning approaches over different sparsity levels in iterative pruning.
Sparsity Magnitude Random DTC
0.1016 0.9638 0.9637 0.9663
0.1914 0.9675 0.9669 0.9687
0.3477 0.9735 0.9723 0.9728
0.4141 0.9745 0.9736 0.9735
0.5254 0.9740 0.9747 0.9732
0.6172 0.9730 0.9748 0.9732
0.7520 0.9718 0.9719 0.9705
0.8594 0.9560 0.9588 0.9609
0.9121 0.9324 0.9399 0.9463
0.9219 0.9234 0.9341 0.9423
0.9297 0.9144 0.9292 0.9371
0.9512 0.8703 0.9157 0.9153
0.9707 0.8293 0.8834 0.8717
0.9785 0.8138 0.8673 0.8500
0.9824 0.8060 0.8589 0.8391
6.1 Learning performance over compression cycles
We record the test accuracy of the reduced architecture after each train-compress cycle for different pruning approaches: magnitude only, random and DTC based magnitude pruning. As shown in Table 1, Figure 1 and Figure 2. DTC and random pruning perform superior. We observe that DTC shows a good performance even against random pruning after 85 percent sparsity. However, random pruning is shown remarkable resilience. But we have clearly shown that DTC outperforms magnitude pruning.
For very high sparsity levels after 95 percent, randomness shows a bit better performance. This isn’t related to network initialization, a fixed Xavier was used making results on the table fully repeatable. While compression process is robust to initialization that performance fluctuations are negligible, rather this is due to fact that after 95 percent compression network remain with handful of neurons in the hidden layer. Depending on requirements very good performance is also retained even after 95 percent of the neurons are pruned.
Figure 5: Next layer Gibbs entropy of computed measurements within DTC over compression process.
Total weight rays in Figure 3 are shown to be more pronounced for very high sparsity levels. This signifies, attenuation of learned weight as model gets much smaller to compensate the performance loss. This is a striking result as DTC sparsify the weights via inverse compressed sensing.
6.2 Evidence for the proposition
We also reported Gibbs entropy over train-compress iterative pruning in Figures 4 and 5. As complexity reduces, Gibbs entropy over measurement vectors also decreases. Pearson correlations between these and performance reduction are computed for DTC and random pruning. They read 0.9174 and 0.9412 respectively, empirically supporting our proposition.This result demonstrated a strong empirical support for the Theorem 1, Gibbs-randomness compression proposition. While both DTC and random pruning have obviously random mechanism the Gibbs-compression proposition requires. As random pruning correlations is much higher, this strengthen the proposition’s statement. Moreover comonotonic behavior is also verified by the visual inspection between Figure 1 and 3 versus Figure 4 and 5.
6.3 Limitations and open improvements
We have developed foundations in developing unified view of distinct ideas of utilization of compressed sensing, dual-tomographic compression and the proposition that randomness and compression are highly correlated. Our experiments shown first promising steps supporting evidence for the proposition and utility of the DTC approach. We understand that more in-depth experimentation are required in advancing and supporting further these proposed methods and the proposition given in Theorem 1.
Further study on the effect of frequency of dual tomographic procedure during train-compress iterations, more optimizations on the penalty term of inverse compressed sensing procedure, the effect of binning on the Gibbs entropy on the measurement vectors and demonstrations with different architectures appears to be good avenues to explore for expanded future studies that requires much larger compute resources for advancing the approach with more diverse datasets.
7 Conclusion
There are multiple original contributions of the work appear to be quite stimulating for both theory and practice. 1) The formulation of an inverse compressed sensing framework, where measurements are generated rather than being a primary optimization process, 2) a new train-compress iterative pruning approach at neuronal level (layers reduced in one go, rather than pruning connections) for deep learning, via new technique called Dual-Tomographic Compression (DTC), is shown as a promising new approach for immensely energy efficient deep learning, 3) demonstration of accelerated lottery tickets hypothesis, that we can find sub-networks much faster during train-compress cycle rather that fully train and find the subnetwork, 4) we have formed mathematically more explicit relationship between directed randomness and complexity reduction compared to algorithmic randomness framework of SKC via a simple theorem, and demonstrated this empirically on our train-compress framework on the canonical deep learning task of classifying handwritten images as a promising results.
Acknowledgment
We are grateful to Y.Süzen for her kind support of this work and her role in securing private funding. Especially her encouragement to make the prototype codes public domain in order to contribute to energy-efficient AI for a better world.
Author is also grateful for constructive comments from the Neural Networks community, strengthen the paper’s technical case and presentation.
Supplement
Two Jupyter notebooks are provided for prototype implementations and reproducing the results presented in this work, named dtcp_analysis.ipynbdtcp\_analysis.ipynb and dtcp_generate_data.ipynbdtcp\_generate\_data.ipynb